# Quantum diffeomorphisms cannot make indefinite causal order definite

### v2

@@ Abstract
Now, $s$ can be determined by measuring the control, together with the spacetime metric, in the basis $\{\mathinner{|{\phi_{\pm}=}\rangle}\frac{1}{\sqrt{2}}(\mathinner{|{1}\rangle}\mathinner{|{g^{(1)}}\rangle}\pm\mathinner{|{2}\rangle}\mathinner{|{g^{(2)}}\rangle})\}$. By post-selecting on the outcome $\mathinner{|{\phi_{+}}\rangle}$, the test particle is left in a superposition of states corresponding to opposite causal orders,
@@
The same procedure can be applied in the case of a definite causal order. It is easy to see that in these cases, the final state before post-selection becomes $(\alpha\mathinner{|{1}\rangle}\mathinner{|{g^{(1)}}\rangle}+\beta\mathinner{|{2}\rangle}\mathinner{|{g^{(2)}}\rangle})\mathinner{|{b_{f}}\rangle}|s=\pm 1\rangle$. In the appendix, we provide details on a measurement scheme to determine the state of the test particle and distinguish between definite and coherently or incoherently superposed causal order.
@@
Discussion — We showed that the proposed notion of ICO qualifies as an observable in both the general relativistic and quantum mechanical sense, and that this observable describes both the optical and the gravitational quantum switch. In particular, we have seen that whether causal order is definite or indefinite is invariant under quantum coordinate transformations. On the other hand, whether ICO is due to delocalized events or a superposition of spacetimes is a matter of choice of quantum coordinates. Thus, we argue that the optical quantum switch can be seen as a realization of indefinite causal order rather than a simulation. Note that this implies that a realization of ICO does not suffice to show that spacetime has been set in a superposition. Moreover, we have seen that there is no quantum coordinate system in which (i) both events ${\mathcal{E}_{A}}$ and ${\mathcal{E}_{B}}$ are localized, (ii) the spacetime metric is definite, and (iii) the causal order between these events is indefinite (cf. [15]).
@@
The present work connects scenarios with localized events on an indefinite spacetime background with delocalized events on definite spacetime. This idea has been prevalent in the community for years [13, 21, 38, 39, 29, 30, 15, 12] but is made explicit here. We are positive that emphasizing this connection will facilitate the discussion between the general relativity and quantum information communities.
@@
What remains to be done is to cast the quantum-controlled diffeomorphisms employed here in the framework of the quantum reference frame formalism. This requires identifying the explicit form of the unitary representations $\mathcal{U}_{1,2}$ and the corresponding Hilbert space structure for the spacetime metric and the quantum coordinates. A promising route to constructing such explicit QRF transformations currently explored is to extend the frameworks of [38, 39, 29].
@@
The present findings are in line with several other works on ICO. First, we can view them in light of recent results in the process matrix formalism. It was shown in [46] that the causal order between operations is always preserved under continuous and reversible transformations (see also [47, 48] for further discussion). This result is derived in an abstract framework of Hilbert spaces in the absence of any spacetime structure and has not yet been extended to a field theoretic realm. However, similarities between the general structure of the transformations in this work and diffeomorphisms hint at a deeper connection, which could explain why causal order is preserved in both frameworks. Second, different notions of ICO have also been analyzed and compared in [15]. The authors connect information-theoretic notions of ICO to relativistic causality and prove several no-go results in the context of localized events, definite spacetime, and classical reference frames. We have already seen consistency with one of their results in the aforementioned no-go-theorem. It would be interesting to investigate to what extent these results extend further to the realm considered here of delocalized events, indefinite metrics and quantum coordinate systems.
@@ Acknowledgements.
MC acknowledges several useful preliminary discussions with Giulio Chiribella and Some Sankar Bhattacharya. We acknowledge financial support by the Austrian Science Fund (FWF) through BeyondC (F7103-N48). This publication was made possible through the support of the ID 61466 and ID 62312 grants from the John Templeton Foundation, as part of The Quantum Information Structure of Spacetime (QISS) Project (qiss.fr). The opinions expressed in this publication are those of the authors and do not necessarily reflect the views of the John Templeton Foundation.
Appendix
@@ Appendix A Mapping between different descriptions of the quantum switch
In this appendix, we take a closer look at the concrete mappings from a description of the quantum switch in which ICO is due to delocalized events to one in which ICO is due to a superposition of spacetime metrics and vice versa.
@@ A.1 Delocalized events to superposition of spacetime metrics
while smoothly connecting to the identity outside of a compact region containing these points. In this way, the events are localized, that is, they are assigned the same respective point on the manifold $\mathcal{M}$, while the initial and final points $p_{\gamma_{0},i}$ and $p_{\gamma_{0},f}$ of the worldline $\gamma_{0}$ remain localized. Let us briefly comment on a common misunderstanding: one might worry that no diffeomorphism can map two different points onto the same one since diffeomorphisms are one-to-one maps. This reasoning does not apply to the present case since we are taking points $p_{{\mathcal{E}_{A}},{\mathcal{E}_{B}}}^{(1)}\in\mathcal{M}_{1}$ and $p_{{\mathcal{E}_{A}},{\mathcal{E}_{B}}}^{(2)}\in\mathcal{M}_{2}$ from two different manifolds $\mathcal{M}_{1}$ and $\mathcal{M}_{2}$ onto the same point $p_{{\mathcal{E}_{A}},{\mathcal{E}_{B}}}\in\mathcal{M}$.
@@
leading to a superposition of metric field values at the points $p_{\mathcal{E}_{A}}$ and $p_{\mathcal{E}_{B}}$, respectively. Note that the metrics $g^{(1)}$ and $g^{(2)}$ are diffeomorphic, namely, they are related by $g^{(2)}={(\phi^{(1)}\circ(\phi^{(2)})^{-1})}_{\ast}g^{(1)}$. Nevertheless, because the metric field at a particular spacetime point is not a diffeomorphism-invariant quantity, it ends up in a superposition of values at this particular point. In fact, the metric field can even be in a superposition of different values along the entire worldline $\gamma_{0}$ between the two events. Diffeomorphism-invariant global geometric properties, on the other hand, are unaffected by this transformation. The two metrics in superposition still both describe a flat spacetime. However, the overall configurations, including both the metric and the worldlines, across the two branches of the superposition are not related by a diffeomorphism. In particular, the diffeomorphism-invariant quantity $s$, that is, the causal order in each branch, is different, since we started with the causal order in a superposition. This fact remains unchanged. If we had started with a situation with definite causal order in the original setup, then no transformation could have made ICO emerge.
@@ A.2 Superposition of spacetime metrics to delocalized events
Thus, the events ${\mathcal{E}_{A}}$ and ${\mathcal{E}_{B}}$ are delocalized. Again, because it is a diffeomorphism-invariant quantity, the causal order remains indefinite. We have thus mapped a situation with ICO due to a superposition of spacetime metrics to a situation with ICO due to delocalized events. Note, importantly, that we have only rendered the metric flat and definite along the worldline $\gamma_{0}$. While this is the only domain relevant for causal order defined through differences of proper time, $\tau_{\gamma}=\frac{1}{c}\int_{p_{\gamma,i}}^{p_{\gamma,f}}\sqrt{-g_{\mu\nu}dx^{\mu}dx^{\nu}}$, the metric generally remains in superposition of values at any point that does not belong to $\gamma_{0}$. Global and diffeomorphism-invariant geometric quantities will not be affected by this transformation.
Remark: In light of more recent work by some of us [49], a more explicit way of modeling the maps between manifolds in superposition would be to retain the distinction between manifolds $\mathcal{M}_{1}$ and $\mathcal{M}_{2}$ even after the application of the diffeomorphism and instead introduce a “comparison map” $C_{\chi}:\mathcal{M}_{1}\to\mathcal{M}_{2}$, constructed from physical fields $\chi$ on the manifolds, that allows us to relate points in one branch of the superposition to points in the other branch and thus define the localization of events. The present results fit into this more general framework by acknowledging that we are implicitly using the identity to compare across the two branches of the superposition (see also [49, p. 22]).
@@ Appendix B An operational encoding of the causal order
In this appendix, we discuss in further detail the operational encoding and verification of indefinite causal order. The protocol proposed in the section Operational encoding as quantum mechanical observable of the main text is illustrated in Fig. 3 below.
Figure 3: In order to encode the causal order operationally, we consider a particle with an internal spin degree of freedom moving along $\gamma_{0}$, and two agents in laboratories with worldlines $\gamma_{A}$ and $\gamma_{B}$. These are placed in a superposition of causal orders between events $\mathcal{E}_{A}$ and $\mathcal{E}_{B}$. The setup is chosen such that the particle always enters the first laboratory at proper time $\tau_{1}^{\ast}$ while it crosses the second laboratory at $\tau_{2}^{\ast}$. Through a careful choice of these proper times, we ensure that the agents can perform non-disturbing measurements of the proper time of the particle whenever it enters their laboratory and thus encode the causal order in a memory register.

### v3

@@ Abstract
Now, $s$ can be determined by measuring the control, together with the spacetime metric, in the basis $\{\mathinner{|{\phi_{\pm}}\rangle}=\frac{1}{\sqrt{2}}(\mathinner{|{1}\rangle}\mathinner{|{g^{(1)}}\rangle}\pm\mathinner{|{2}\rangle}\mathinner{|{g^{(2)}}\rangle})\}$. By post-selecting on the outcome $\mathinner{|{\phi_{+}}\rangle}$, the test particle is left in a superposition of states corresponding to opposite causal orders,
@@
The same procedure can be applied in the case of a definite causal order. It is easy to see that in these cases, the final state before post-selection becomes $(\alpha\mathinner{|{1}\rangle}\mathinner{|{g^{(1)}}\rangle}+\beta\mathinner{|{2}\rangle}\mathinner{|{g^{(2)}}\rangle})\mathinner{|{b_{f}}\rangle}|s=\pm 1\rangle$. In Supplemental Material B, we provide details on a measurement scheme to determine the state of the test particle and distinguish between definite and coherently or incoherently superposed causal order.
@@
Discussion — We proposed a notion of ICO between operationally defined events. One key conceptual contribution is the extension of the notion of an event to superpositions of spacetimes. The advantage of this notion is that it is independent of the choice of classical and quantum coordinates and thus qualifies as an observable in the general relativistic sense. We further showed that the proposed notion of ICO constitutes an observable in the quantum mechanical sense, and that this observable describes both the optical and the gravitational quantum switch. In particular, we have seen that whether causal order is definite or indefinite is invariant under quantum coordinate transformations. On the other hand, whether ICO is due to delocalized events or a superposition of spacetimes is a matter of choice of quantum coordinates. Thus, we argue that the optical quantum switch can be seen as a realization of indefinite causal order rather than a simulation. Note that this implies that a realization of ICO does not suffice to show that spacetime has been set in a superposition. Moreover, we have seen that there is no quantum coordinate system in which (i) both events $\mathcal{E}_{A}$} and $\mathcal{E}_{B}$} are localized, (ii) the spacetime metric is definite, and (iii) the causal order between these events is indefinite (cf. [24]).
@@
The present work connects scenarios with localized events on an indefinite spacetime background with delocalized events on definite spacetime. This idea has been prevalent in the community for years [22, 30, 39, 45, 20, 38, 24, 14] but is made explicit here. We are positive that emphasizing this connection will facilitate the discussion between the general relativity and quantum information communities and shed light on the interplay between indefinite causality and the quantum nature of spacetime. This might also be relevant to the ongoing research efforts regarding potential tests of quantum features of gravity in the low energy regime (e.g. [15, 16, 17, 18, 19, 4, 20, 21]).
@@
What remains to be done is to cast the quantum-controlled diffeomorphisms employed here in the framework of the quantum reference frame formalism. This requires identifying the explicit form of the unitary representations $\mathcal{U}_{1,2}$ and the corresponding Hilbert space structure for the spacetime metric and the quantum coordinates. A promising route to constructing such explicit QRF transformations currently explored is to extend the frameworks of [39, 45, 20]. Another interesting direction for future research is to extend the notion of ICO to more general scenarios. This includes finding an appropriate definition of event when given worldlines with multiple crossings or even different numbers of crossings across the branches of the superposition, as well as generalizing the notion to quantum gravitational scenarios beyond a superposition of semi-classical spacetimes. In particular, an important open question is concerned with the extension of the notion of event to genuine quantum fields. In this context, it would further be interesting to explore the relation between ICO and micro-causality in algebraic quantum field theory (cf. [25], [51, Ch. 3], [52, Ch. 3]).
@@
The present findings are in line with several other works on ICO. First, we can view them in light of recent results in the process matrix formalism. It was shown in [53] that the causal order between operations is always preserved under continuous and reversible transformations (see also [54, 55] for further discussion). This result is derived in an abstract framework of Hilbert spaces in the absence of any spacetime structure and has not yet been extended to a field theoretic realm. However, similarities between the general structure of the transformations in this work and diffeomorphisms hint at a deeper connection, which could explain why causal order is preserved in both frameworks. Second, different notions of ICO have also been analyzed and compared in [24]. The authors connect information-theoretic notions of ICO to relativistic causality and prove several no-go results in the context of localized events, definite spacetime, and classical reference frames. We have already seen consistency with one of their results in the aforementioned no-go-theorem. It would be interesting to investigate to what extent these results extend further to the realm considered here of delocalized events, indefinite metrics, and quantum coordinate systems.
@@ Acknowledgements.
MC acknowledges several useful preliminary discussions with Giulio Chiribella and Some Sankar Bhattacharya. VK acknowledges support through a DOC Fellowship of the Austrian Academy of Sciences. This research was funded in whole or in part by the Austrian Science Fund (FWF) [10.55776/F71]. This publication was made possible through the support of the ID 61466 and ID 62312 grants from the John Templeton Foundation, as part of The Quantum Information Structure of Spacetime (QISS) Project (qiss.fr). The opinions expressed in this publication are those of the authors and do not necessarily reflect the views of the John Templeton Foundation.
Supplemental Material
@@ Appendix A Mapping between different descriptions of the quantum switch
In this Supplemental Material, we take a closer look at the concrete mappings from a description of the quantum switch in which ICO is due to delocalized events to one in which ICO is due to a superposition of spacetime metrics and vice versa.
@@ A.1 Delocalized events to superposition of spacetime metrics
$\phi^{(1)}(p_{{\mathcal{E}_{B}}}^{(1)})$ $=\phi^{(2)}(p_{{\mathcal{E}_{B}}}^{(2)})\equiv p_{{\mathcal{E}_{B}}},$ (14)
while smoothly connecting to the identity outside of a compact region containing these points. In this way, the events are localized, that is, they are assigned the same respective point on the manifold $\mathcal{M}$, while the initial and final points $p_{\gamma_{0},i}$ and $p_{\gamma_{0},f}$ of the worldline $\gamma_{0}$ remain localized. Note that the idea that one can find coordinates such that the events $\mathcal{E}_{A}$} and $\mathcal{E}_{B}$} – in this case defined by the application of operations in $A$ and $B$’s laboratory – become localized was also discussed in [13]. Let us also briefly comment on a common misunderstanding: one might worry that no diffeomorphism can map two different points onto the same one since diffeomorphisms are one-to-one maps. This reasoning does not apply to the present case since we are taking points $p_{{\mathcal{E}_{A}},{\mathcal{E}_{B}}}^{(1)}\in\mathcal{M}_{1}$ and $p_{{\mathcal{E}_{A}},{\mathcal{E}_{B}}}^{(2)}\in\mathcal{M}_{2}$ from two different manifolds $\mathcal{M}_{1}$ and $\mathcal{M}_{2}$ onto the same point $p_{{\mathcal{E}_{A}},{\mathcal{E}_{B}}}\in\mathcal{M}$.
@@
leading to a superposition of metric field values at the points $p_{\mathcal{E}_{A}}$ and $p_{\mathcal{E}_{B}}$, respectively. Note that the metrics $g^{(1)}$ and $g^{(2)}$ are diffeomorphic, namely, they are related by $g^{(2)}={(\phi^{(1)}\circ(\phi^{(2)})^{-1})}_{\ast}(g^{(1)})$. Nevertheless, because the metric field at a particular spacetime point is not a diffeomorphism-invariant quantity, it ends up in a superposition of values at this particular point. In fact, the metric field can even be in a superposition of different values along the entire worldline $\gamma_{0}$ between the two events. Diffeomorphism-invariant global geometric properties, on the other hand, are unaffected by this transformation. The two metrics in superposition still both describe a flat spacetime. However, the overall configurations, including both the metric and the worldlines, across the two branches of the superposition are not related by a diffeomorphism. In particular, the diffeomorphism-invariant quantity $s$, that is, the causal order in each branch, is different, since we started with the causal order in a superposition. This fact remains unchanged. If we had started with a situation with definite causal order in the original setup, then no transformation could have made ICO emerge.
@@ A.2 Superposition of spacetime metrics to delocalized events
Thus, the events $\mathcal{E}_{A}$} and $\mathcal{E}_{B}$} are delocalized. Again, because it is a diffeomorphism-invariant quantity, the causal order remains indefinite. We have thus mapped a situation with ICO due to a superposition of spacetime metrics to a situation with ICO due to delocalized events. Note, importantly, that we have only rendered the metric flat and definite along the worldline $\gamma_{0}$. While this is the only domain relevant for causal order defined through differences of proper time, $\tau_{\gamma_{0}}=\frac{1}{c}\int_{p_{\gamma_{0},i}}^{p_{\gamma_{0},f}}\sqrt{-g_{\mu\nu}dx^{\mu}dx^{\nu}}$, the metric generally remains in superposition of values at any point that does not belong to $\gamma_{0}$. Global and diffeomorphism-invariant geometric quantities will not be affected by this transformation.
Remark: In light of more recent work by some of us [56], a more explicit way of modeling the maps between manifolds in superposition would be to retain the distinction between manifolds $\mathcal{M}_{1}$ and $\mathcal{M}_{2}$ even after the application of the diffeomorphism and instead introduce a “comparison map” $C_{\chi}:\mathcal{M}_{1}\to\mathcal{M}_{2}$, constructed from physical fields $\chi$ on the manifolds, that allows us to relate points in one branch of the superposition to points in the other branch and thus define the localization of events. The present results fit into this more general framework by acknowledging that we are implicitly using the identity to compare across the two branches of the superposition (see also [56, p. 22]).
@@ Appendix B An operational encoding of the causal order
In this Supplemental Note, we discuss in further detail the operational encoding and verification of indefinite causal order. The protocol proposed in the section Operational encoding as quantum mechanical observable of the main text is illustrated in Fig. 3 below.
Figure 3: In order to encode the causal order operationally, we consider a particle with an internal spin degree of freedom moving along $\gamma_{0}$, and two agents in laboratories with worldlines $\gamma_{A}$ and $\gamma_{B}$. These are placed in a superposition of causal orders between events $\mathcal{E}_{A}$ and $\mathcal{E}_{B}$. The setup is chosen such that the particle always enters the first laboratory at proper time $\tau_{1}^{\ast}$ while it crosses the second laboratory at $\tau_{2}^{\ast}$. Through a careful choice of these proper times, we ensure that the agents can perform non-disturbing measurements on the particle whenever it enters their laboratory and thus encode the causal order in a memory register.